# Entanglement measures of Majorana bound states

### v1

@@ III.4 Generating a maximally entangled state between MBSs and a QD
As pointed out before, the oscillatory behaviour of the entanglement measures as functions of time imply a possible generation of maximally entangled states. Since the considered systems are composed of three subsystems (qubits), with an initial state of maximally entangled QDs, we focus on achieving a maximally entangled state between MBSs and the right QD. To find a maximally entangled state, it is necessary to have $\eta=\chi$ in Eq. (19), because this conditions places the time-evolved state given by Eq. (18) in the form of two qubit Bell’s states, which are maximally entangled states [41]. Thus, taking this condition into account, we find the required parameters for achieving maximal entanglement. Then, we can choose $t=((2n+1)\pi/{2\Delta})\,\hbar/\lambda_{1}$ so that $\eta$ is purely imaginary; then, by comparing the imaginary parts of $\chi$ and $\eta$, we find that they are equal when $\omega=1-\lambda$, which, using the expression for $\Delta$ below Eq. (3), gives $\Delta=\sqrt{2}|1-\lambda|$. Under these conditions, we obtain $\eta=\chi={1}/{\sqrt{2}}$, which is expected to lead to a maximally entangled state between MBSs and QDs because, as we explained above, the time-evolved state acquires the form of two-qubit Bell states.
@@
Figure 3: Generation of a maximally entangled state between MBSs with a QD. (a,b) Time evolution of concurrence, discord, return probability, and entanglement dynamics, denoted by $C$, $D$, $R_{p}$, and $E_{d}$, with an initial state of maximally entangled QDs for MS and NFS. Here, $\lambda=0.7$ and $\omega=1-\lambda$ in panels (a,b). The red star in (a) shows that the QDs are maximally entangled at the beginning of the evolution $t=0$, whereas the magenta stars show that maximally entanglement between MBSs and the second QD is achieved at later times. (c) Schematic representation of the entangled systems indicated by red and magenta stars in (a), with the thick cyan arrow showing the entangled parts. The NFS does not have this characteristic.
@@
In spite of the seemingly stringent conditions to achieve maximally entanglement during the evolution, in Fig. 3 we show that it is possible at $\omega=1-\lambda$ and $\lambda=0.7$ by plotting the time evolution of $R_{p}$, $E_{d}$, $C$, and $D$ for the MS and NFS setups. Apart from the already seen oscillatory behaviour, there are two points we would like to highlight, which are marked by red and magenta stars in Fig. 3(a) for the MS setup. First, at the beginning of the evolution $t=0$, we have $R_{p}=1$, $C=1$, $D=1$, and $E_{d}=0.25$, a regime that corresponds to the maximally entangled QDs. Second, at times given by $t=\pi/{2\Delta}=3.7\,\hbar/\lambda_{1}$, the entanglement dynamics and return probability become $E_{d}=0.5$ and $R_{p}=0.5$, implying the creation of a maximally entangled state of MBSs and the second QD, see magenta star in Fig. 3(a). This maximal entanglement is further confirmed by getting vanishing concurrence $C=0$ at such times, a situation that displays the monogamy behaviour of entanglement: when MBSs and the right QD are maximally entangled, the right QD cannot be entangled with the left QD, thus leading to $C=0$. Interestingly, the vanishing concurrence of the MS ($C=0$) is accompanied by vanishing discord ($D=0$), which reveals the absence of quantum correlations and that the system transitions into a classical state, an effect only obtained in the MS but not in the NFS, see Fig. 3(a,b). To further visualize the achieved entanglement between MBSs and right QD, in Fig. 3(c) we schematically illustrate the entanglement shift that corresponds to the two times of the evolution indicated by red and magenta stars in Fig. 3(a). This entanglement shift reveals the possibility of starting with two maximally entangled QDs and then generate maximally entanglement between MBSs and right QD during the time evolution. In contrast to the MS system, the entanglement measures for the NFS develop a profile where it is not simple to determine the generation maximal entanglement between MBSs and QDs, Fig. 3(b); their evolution is similar to Fig. 2(b,d,f,h). Therefore, it is possible to induce maximally entangled states between MBSs and QDs, even though for this purpose finite energy MBSs are needed, which, however, are not nonlocal and might not serve the purpose of encoding information nonlocally.
@@ IV Results for a separable initial state
where $\lambda$ characterizes the coupling to the right QD, $\omega$ is the energy of MBSs (normal fermion), $\Delta$, $\Delta_{1}$, and $\Delta_{pm}$ are given below Eqs. (3) and Eqs. (6). It is important to note that all the coefficients for MS setup are finite and hence a complex time evolution is expected. Moreover, contrary to the coefficients of |$\ket{\psi(t)}$ for initially maximally entangled states of the QDs in Eq. (25) for NFS, the first coefficient, $\bar{c}^{\prime}_{1}(t)$, goes to zero. This occurs because the state |$\ket{000}$ is an eigenstate of the system, and the initial state does not contain the configuration |$\ket{000}$. Therefore, it does not contribute to the evolution (see Eq. (17)). The other coefficients are non-zero and show a similar form of dependence on $\lambda$ and $\omega$, as we will observe in the dynamics. Using the coefficients from Eq. (28) and Eq. (29) for MS and NFS, and following the steps to calculate the state dynamical function and quantum correlations given in Section III, we obtain $E_{d}$, $C$, and $D$ and discuss their time evolution next.
IV.1 Discussion of the state dynamics and quantum correlations
@@
In the case of symmetric couplings ($\lambda=1$), the immediate observation is that almost all the quantities oscillate with time, with different patterns at $\omega=0$ and $\omega\neq 0$, which is different to what we observed in the previous section for initial maximally entangled QDs. For the MS case in Fig. 4(a), the quantities $R_{p}$, $E_{d}$, and $D$ oscillate, while $C=0$, implying that the completely nonlocal zero-energy MBSs cannot generate entanglement in this case. For the entanglement dynamics and return probability, we obtain $E_{d}=\sin^{2}t$ and $R_{p}=\cos^{2}t$, which, at $t=(\pi/2)\,\hbar/\lambda_{1}$, become $E_{d}=0.5$ and $R_{p}=0$, as depicted by black and green curves in Fig. 4(a). By obtaining $c_{1}(t)=1$, we conclude that at those points in time, the state |$\ket{000}$ is present. In relation to the discord $D$ in Fig. 4(a), it exhibits homogeneous oscillations as a function of time, with period equal to $T=(\pi/4)\,\hbar/\lambda_{1}$, acquiring vanishing values at points where $C$ vanishes, a phenomenon we have seen to occur only for MS in Figs. 2 and Fig. 3 of the previous section. The times at which the discord completes a period ($\pi/4$) are special because the eigenvalues of the marginal entropies are $\{0,1\}$ at $\theta=\pi/4$ and $\phi=\pi/4$, which then gives vanishing conditional entropy when obtaining the discord in Eq. 13, see also Eq. (36) and Eq. (37) in Appendix A.1. Moreover, the composite entropy and the single QD entropy are the same and equal to unity, thus the total algebraic sum goes to zero, see details in Eq. (37) in Appendix A.1. When analyzing the time $t=(\pi/2)\,\hbar/\lambda_{1}$, we have $c_{1}(t)=1$ in Eq. (28) which gives a pure state description, where all entropies go to zero. For a finite overlap ($\lambda=1,\omega=0.7$), plotted in Fig. 4(c), an extra frequency in oscillation is introduced because of $\omega$. The entanglement is generated between QDs with fast and slow frequencies. In this case, $C$ and $D$ reach unity simultaneously, implying that maximum quantum correlation can be created by MBSs at finite energy splitting [62]. We conclude that, while quantum correlations can be generated with zero-energy MBSs but they remain small, entanglement generation is possible only with a finite energy splitting of MBSs.
@@
For asymmetric couplings at $\lambda=0.7$, the situation is slightly different but with some similarities, see Fig. 4(e,g) for $\omega=0$ and $\omega\neq 0$ in the MS setup. First of all, $R_{p}$, $E_{d}$, and $D$ oscillate with time, see Fig. 4(e,g). As for the symmetric case, $C=0$ throughout the evolution at $\omega=0$ [Fig. 4(e)], meaning that entanglement can not be generated. This is because the eigenenergies in Eqs. (4) are given by $\Delta=|\lambda-1|$, which simplifies the coefficients in Eq. (28) as $c_{1(4)}(t)=[\cos{(\lambda-1)t}\mp\cos{(\lambda+1)t}]/2$, $c_{2(3)}(t)=\pm i[\sin{(\lambda+1)t}\mp\sin{(\lambda-1)t}]/2$. This gives $|c_{1}(t)||c_{4}(t)|=|c_{2}(t)||c_{3}(t)|=|\sin^{2}{(\lambda-1)t}-\sin^{2}{(\lambda+1)t}|$. Therefore, the concurrence $C$, as defined in Eq. (26), goes to zero throughout the evolution, see Fig. 4(e). However, at $\omega\neq 0$, $C$ takes finite values and develops an oscillatory profile, see Fig. 4(g). In relation to the discord, it develops oscillations with different periodicities, acquiring a beating profile that oscillates faster for finite frequencies, as seen in Fig. 4(e,g). We note that the complex behaviour of $D$ stems from the fact that it is determined by different entropies, see Appendix A.1. When it comes to the entanglement dynamics in asymmetric couplings ($\lambda\neq 1$) with $\omega=0$, we find that $E_{d}=\sin^{2}{t}/2$ which is independent of $\lambda$ and has a constant period of $\pi$ as seen Fig. 4(e). In contrast, in the same regime, we obtain the return probability to be $R_{p}=\cos^{2}{\lambda t}\cos^{2}t$, which clearly depends on two periods, $\pi/\lambda$ and $\pi$. Therefore, the system takes a long time to return to the initial state, beyond the scope of Fig. 4(e). For clarification, we note that the first maxima of $R_{p}$ happens at $t=(10\pi)\,\hbar/\lambda_{1}$. Moving into the asymmetric couplings ($\lambda\neq 1$) and finite $\omega\neq 0$, as shown in Fig. 4(g), the quantities $R_{p}$, $E_{d}$, $C$, and $D$ exhibit an oscillatory dynamics without showing any special entanglement characteristics, unlike their counterparts for a maximally entangled initial state discussed in Section III.
@@
In relation to the entanglement measures for the NFS system, shown in Fig. 4(b,d,f,h), they exhibit a beaviour that is roughly similar either at finite or zero energies and at symmetric or asymmetric couplings. Here, the concurrence we obtain to be equal to $C=||\bar{c}^{\prime}_{2}(t)||\bar{c}^{\prime}_{3}(t)||$, which goes to zero when either of the coefficients approaches zero. The time at which $C=0$ is calculated as $t=(n\pi/{\sqrt{1+\lambda^{2}+\omega^{2}}})\,\hbar/\lambda_{1}$, where n is an integer. At these points in time, the NFS acquires the state |$\ket{110}$, which is a zero concurrence state ($C=0$). This can be verified from the results in Fig. 4(b,d,f,h), where zero concurrence occurs between $t=2\,\hbar/\lambda_{1}$ and $t=3\,\hbar/\lambda_{1}$, depending on the particular value of $\lambda$ and $\omega$ during the evolution.
@@ IV.2 Generating maximally entangled QDs induced by MBSs
We have seen in the previous subsection that the time-evolution in the MS setup is sensitive to variations of $\omega$. Of particular interest in this part is the behaviour of the quantum correlations at $\omega\sim 0$, because this regime has well localized MBSs in the MS. Here we exploit the sensitivity of discord $D$ at $\omega\sim 0$ to generate a maximally entangled state between QDs. The time evolution of $D$ at $\omega=0.001$ and asymmetric coupling $\lambda=0.7$ is presented in Fig. 5(a,b) for the MS setup for long and short times, with Fig. 5(b) taken within the green box Fig. 5(a). We contrast this behavior with the time evolution for the NFS at short times. The first observation is that $D$ exhibits rapid oscillations with an initial small amplitude of $0.1$ at short times [Fig. 5(b)]. As time progresses, $D$ increases and eventually reaches unity in the long-term evolution, see Fig. 5(a). Having $D=1$ means that a maximally entangled stated of QDs is created, achieved only at a small energy splitting of MBSs. In contrast to the MS, the discord for the NFS system in Fig. 5(c) develops homogeneous oscillations at short times whose period and amplitudes do not change over time. We have verified that this behaviour in the NFS also remains in the long time dynamics, making it very different to the MS.
V Conclusions

### v2

@@ III.4 Generating a maximally entangled state between MBSs and a QD
As pointed out before, the oscillatory behaviour of the entanglement measures as functions of time implies a possible generation of maximally entangled states. Since the considered systems are composed of three subsystems (qubits), with an initial state of maximally entangled QDs, we focus on achieving a maximally entangled state between MBSs and the right QD. To find a maximally entangled state, it is necessary to have $\eta=\chi$ in Eq. (20), because this conditions places the time-evolved state given by Eq. (19) in the form of two qubit Bell’s states, which are maximally entangled states [71]. Thus, taking this condition into account, we find the required parameters for achieving maximal entanglement. Then, we can choose $t=((2n+1)\pi/{2\Delta})\,\hbar/\lambda_{1}$ so that $\eta$ is purely imaginary; then, by comparing the imaginary parts of $\chi$ and $\eta$, we find that they are equal when $\omega=1-\lambda$, which, using the expression for $\Delta$ below Eq. (3), gives $\Delta=\sqrt{2}|1-\lambda|$. Under these conditions, we obtain $\eta=\chi={1}/{\sqrt{2}}$, which is expected to lead to a maximally entangled state between MBSs and QDs because, as we explained above, the time-evolved state acquires the form of two-qubit Bell states.
@@
Figure 3: Generation of a maximally entangled state between MBSs with a QD. (a,b) Time evolution of concurrence, discord, return probability, and entanglement dynamics, denoted by $C$, $D$, $R_{p}$, and $E_{d}$, with an initial state of maximally entangled QDs for MS and NFS. Here, $\lambda=0.7$ and $\omega=1-\lambda$ in panels (a,b). The red star in (a) shows that the QDs are maximally entangled at the beginning of the evolution $t=0$, whereas the magenta stars show that maximal entanglement between MBSs and the second QD is achieved at later times. (c) Schematic representation of the entangled systems indicated by red and magenta stars in (a), with the thick cyan arrow showing the entangled parts. The NFS does not have this characteristic.
@@
In spite of the seemingly stringent conditions to achieve maximal entanglement during the evolution, in Fig. 3 we show that it is possible at $\omega=1-\lambda$ and $\lambda=0.7$ by plotting the time evolution of $R_{p}$, $E_{d}$, $C$, and $D$ for the MS and NFS setups. Apart from the already seen oscillatory behaviour, there are two points we would like to highlight, which are marked by red and magenta stars in Fig. 3(a) for the MS setup. First, at the beginning of the evolution $t=0$, we have $R_{p}=1$, $C=1$, $D=1$, and $E_{d}=0.25$, a regime that corresponds to the maximally entangled QDs. Second, at times given by $t=\pi/{2\Delta}=3.7\,\hbar/\lambda_{1}$, the entanglement dynamics and return probability become $E_{d}=0.5$ and $R_{p}=0.5$, implying the creation of a maximally entangled state of MBSs and the second QD, see magenta star in Fig. 3(a). This maximal entanglement is further confirmed by getting vanishing concurrence $C=0$ at such times, a situation that displays the monogamy behaviour of entanglement: when MBSs and the right QD are maximally entangled, the right QD cannot be entangled with the left QD, thus leading to $C=0$. Interestingly, the vanishing concurrence of the MS ($C=0$) is accompanied by vanishing discord ($D=0$), which reveals the absence of quantum correlations and that the system transitions into a classical state, an effect only obtained in the MS but not in the NFS; see Fig. 3(a,b).
The key difference in achieving the classical state in the MS, as opposed to the NFS, originates from the energy contribution arising from $\omega\neq 0$. This drives the dynamics in a special way that the time-evolved state can be expressed in the Bell’s states form with the same coefficients for odd and even parity of the first QD, see Eq. (19). In contrast, the energy contribution in the NFS due to the normal fermion creates an arbitrarily evolving state, where each configuration has different amplitudes, as seen in Eq. (25). These distinct time-evolved states for the two systems lead to differences in the evolution of quantum correlations over time. In the MS, the classical state is achieved at different times when the conditional entropy, which captures nonclassical correlations, goes to zero, coinciding with the vanishing of entanglement. Meanwhile, in the NFS, the conditional entropy remains significant at those same moments, quantifying the nonclassical correlations in the system.
To further visualize the achieved entanglement between MBSs and right QD, in Fig. 3(c) we schematically illustrate the entanglement shift that corresponds to the two times of the evolution indicated by red and magenta stars in Fig. 3(a). This entanglement shift reveals the possibility of starting with two maximally entangled QDs and then generating maximal entanglement between MBSs and the right QD during the time evolution. In contrast to the MS system, the entanglement measures for the NFS exhibit a profile where the generation of maximal entanglement between the local fermion and the QDs cannot be determined, as shown in Fig. 3(b). Their evolution closely resembles that depicted in Fig. 2(b, d, f, h). Thus, we conclude that maximal entanglement between MBSs and QDs can be induced only in the MS. However, this requires finite-energy MBSs, which are not fully nonlocal and therefore may not be ideal for encoding information nonlocally.
@@ IV Results for a separable initial state
where $\lambda$ characterizes the coupling to the right QD, $\omega$ is the energy of MBSs (normal fermion), $\Delta$, $\Delta_{1}$, and $\Delta_{\pm}$ are given below Eqs. (3) and Eqs. (6). It is important to note that all the coefficients for MS setup are finite, and hence a complex time evolution is expected. Moreover, contrary to the coefficients of |$\ket{\psi(t)}$ for initially maximally entangled states of the QDs in Eq. (26) for NFS, the first coefficient, $\bar{c}^{\prime}_{1}(t)$, goes to zero. This occurs because the state |$\ket{000}$ is an eigenstate of the system, and the initial state does not contain the configuration |$\ket{000}$. Therefore, it does not contribute to the evolution (see Eq. (18)). The other coefficients are non-zero and show a similar form of dependence on $\lambda$ and $\omega$, as we will observe in the dynamics. Using the coefficients from Eq. (29) and Eq. (30) for MS and NFS, and following the steps to calculate the state dynamical function and quantum correlations given in Section III, we obtain $E_{d}$, $C$, and $D$ and discuss their time evolution next.
IV.1 Discussion of state dynamics and quantum correlations
@@
In the case of symmetric couplings ($\lambda=1$), the immediate observation is that almost all the quantities oscillate with time, with different patterns at $\omega=0$ and $\omega\neq 0$, which is different to what we observed in the previous section for initial maximally entangled QDs. For the MS case in Fig. 4(a), the quantities $R_{p}$, $E_{d}$, and $D$ oscillate, while $C=0$, implying that the completely nonlocal zero-energy MBSs cannot generate entanglement in this case. For the entanglement dynamics and return probability, we obtain $E_{d}=\sin^{2}(t)/2$ and $R_{p}=\cos^{2}t$, which, at $t=(\pi/2)\,\hbar/\lambda_{1}$, become $E_{d}=0.5$ and $R_{p}=0$, as depicted by black and green curves in Fig. 4(a). By obtaining $c_{1}(t)=1$, we conclude that at those points in time, the state |$\ket{000}$ is present. In relation to the discord $D$ in Fig. 4(a), it exhibits homogeneous oscillations as a function of time, with period equal to $T=(\pi/4)\,\hbar/\lambda_{1}$, acquiring vanishing values at points where $C$ vanishes, a phenomenon we have seen to occur only for MS in Figs. 2 and Fig. 3 of the previous section. The times at which the discord completes a period ($\pi/4$) are special because the eigenvalues of the marginal entropies are $\{0,1\}$ at $\theta=\pi/4$ and $\phi=\pi/4$, which then gives vanishing conditional entropy when obtaining the discord in Eq. 13, see also Eq. (37) and Eq. (38) in Appendix A.1. Moreover, the composite entropy and the single QD entropy are the same and equal to unity, thus the total algebraic sum goes to zero; see details in Eq. (38) in Appendix A.1. When analyzing the time $t=(\pi/2)\,\hbar/\lambda_{1}$, we have $c_{1}(t)=1$ in Eq. (29) which gives a pure state description, where all entropies go to zero. For a finite overlap ($\lambda=1,\omega=0.7$), plotted in Fig. 4(c), an extra frequency in oscillation is introduced because of $\omega$. The entanglement is generated between QDs with fast and slow frequencies. In this case, $C$ and $D$ reach unity simultaneously, implying that maximum quantum correlation can be created by MBSs at finite energy splitting [94]. We conclude that, while quantum correlations can be generated with zero-energy MBSs but they remain small, entanglement generation is possible only with a finite energy splitting of MBSs.
@@
For asymmetric couplings at $\lambda=0.7$, the situation is slightly different but with some similarities; see Fig. 4(e,g) for $\omega=0$ and $\omega\neq 0$ in the MS setup. First of all, $R_{p}$, $E_{d}$, and $D$ oscillate with time; see Fig. 4(e,g). As for the symmetric case, $C=0$ throughout the evolution at $\omega=0$ [Fig. 4(e)], meaning that entanglement can not be generated. This is because the eigenenergies in Eqs. (4) are given by $\Delta=|\lambda-1|$, which simplifies the coefficients in Eq. (29) as $c_{1(4)}(t)=[\cos{(\lambda-1)t}\mp\cos{(\lambda+1)t}]/2$, $c_{2(3)}(t)=\pm i[\sin{(\lambda+1)t}\mp\sin{(\lambda-1)t}]/2$. This gives $|c_{1}(t)||c_{4}(t)|=|c_{2}(t)||c_{3}(t)|=|\sin^{2}{(\lambda-1)t}-\sin^{2}{(\lambda+1)t}|$. Therefore, the concurrence $C$, as defined in Eq. (27), goes to zero throughout the evolution; see Fig. 4(e). However, at $\omega\neq 0$, $C$ takes finite values and develops an oscillatory profile; see Fig. 4(g). In relation to the discord, it develops oscillations with different periodicities, acquiring a beating profile that oscillates faster for finite frequencies, as seen in Fig. 4(e,g). We note that the complex behaviour of $D$ stems from the fact that it is determined by different entropies; see Appendix A.1. When it comes to the entanglement dynamics in asymmetric couplings ($\lambda\neq 1$) with $\omega=0$, we find that $E_{d}=\sin^{2}(t)/2$ which is independent of $\lambda$ and has a constant period of $\pi$ as seen Fig. 4(e). In contrast, in the same regime, we obtain the return probability to be $R_{p}=\cos^{2}(\lambda t)\cos^{2}(t)$, which clearly depends on two periods, $\pi/\lambda$ and $\pi$. Therefore, the system takes a long time to return to the initial state, beyond the scope of Fig. 4(e). For clarification, we note that the first maxima of $R_{p}$ happens at $t=(10\pi)\,\hbar/\lambda_{1}$. Moving into the asymmetric couplings ($\lambda\neq 1$) and finite $\omega\neq 0$, as shown in Fig. 4(g), the quantities $R_{p}$, $E_{d}$, $C$, and $D$ exhibit oscillatory dynamics without showing any special entanglement characteristics, unlike their counterparts for a maximally entangled initial state discussed in Section III.
@@
In relation to the entanglement measures for the NFS system, shown in Fig. 4(b,d,f,h), they exhibit a beaviour that is roughly similar either at finite or zero energies and at symmetric or asymmetric couplings. Here, the concurrence we obtain is equal to $C=||\bar{c}^{\prime}_{2}(t)||\bar{c}^{\prime}_{3}(t)||$, which goes to zero when either of the coefficients approaches zero. The time at which $C=0$ is calculated as $t=(n\pi/{\sqrt{1+\lambda^{2}+\omega^{2}}})\,\hbar/\lambda_{1}$, where n is an integer. At these points in time, the NFS acquires the state |$\ket{110}$, which is a zero concurrence state ($C=0$). This can be verified from the results in Fig. 4(b,d,f,h), where zero concurrence occurs between $t=2\,\hbar/\lambda_{1}$ and $t=3\,\hbar/\lambda_{1}$, depending on the particular value of $\lambda$ and $\omega$ during the evolution.
@@ IV.2 Generating maximally entangled QDs induced by MBSs
We have seen in the previous subsection that the time-evolution in the MS setup is sensitive to variations of $\omega$. Of particular interest in this part is the behaviour of the quantum correlations at $\omega\sim 0$, because this regime has well localized MBSs in the MS. Here we exploit the sensitivity of discord $D$ at $\omega\sim 0$ to generate a maximally entangled state between QDs. The time evolution of $D$ at $\omega=0.001$ and asymmetric coupling $\lambda=0.7$ is presented in Fig. 5(a,b) for the MS setup for long and short times, with Fig. 5(b) taken within the green box Fig. 5(a). We contrast this behavior with the time evolution for the NFS at short times. The first observation is that $D$ exhibits rapid oscillations with an initial small amplitude of $0.1$ at short times [Fig. 5(b)]. As time progresses, $D$ increases and eventually reaches unity in the long-term evolution, see Fig. 5(a). Having $D=1$ means that a maximally entangled stated of QDs is created, achieved only at a small energy splitting of MBSs. In contrast to the MS, the discord for the NFS system in Fig. 5(c) develops homogeneous oscillations at short times whose period and amplitudes do not change over time. We have verified that this behaviour in the NFS also remains in the long-time dynamics, making it very different to the MS.
V Concluding remarks